OpenStudy (anonymous):

please help@ write an equation of a line in slope intercept form that is paralell to x+4y=6 and passes through the point (-8,5)

OpenStudy (anonymous):

(1) Slope intercept form is y=mx+b (2) Parallel lines all have the same slope, so the slope for this parallel line should have the same slope as x+4y=6 , but a different b (3) If you know what m is in y=mx+b from (2) then you can plug in x,y values to find b 5=m(-8)+b and solve for b. Then you write the equation as y=mx+b with the m you found in (2) and the b you just found.

OpenStudy (anonymous):

slope=-a/b=-1/4 second method by differrential 1+4dy/dx=0 dy/dx=-1/4